OpenStudy (shealah):

Help find the centroid algebraically? G(1,0), H(1,6) and I(10,0)

OpenStudy (shealah):

GI midpoint = J(5,0) GH midpoint = K(1,3)

OpenStudy (shealah):

?

OpenStudy (anonymous):

The centroid is where the medians intersect right?

OpenStudy (shealah):

Yeah

OpenStudy (anonymous):

You can just average the x and y coordinates.

OpenStudy (anonymous):

so\[(\frac{ x _{1}+x _{2}+x _{3} }{ 3}, \frac{ y _{1}+y _{2}+y _{3} }{ 3})\]those are the coordinates of the centroid

OpenStudy (anonymous):

just input the numbers and solve